Suppose a firm has 33 million shares of common stock outstanding at a price of $40 per share. The firm also has 200,000 bonds outstanding with a current price of $1005.2. The outstanding bonds have yield to maturity 10.8%. The firm's common stock beta is 1.2 and the corporate tax rate is 36%. The expected market return is 12% and the T-bill rate is 6%. What is the WACC for this firm? (All answers should be in the format of 0.XXX)

Weight of Equity (3 decimals):

Weight of Debt (3 decimals):

Cost of Equity (3 decimals):

After tax Cost of Debt (3 decimals):

WACC (3 decimals):
Answer #1

Total Equity = 33*40 = $1,320 million

Total Debt = 200,000*1005.20 = $201.04 million

Weight of Equity = 1,320/(1,320 + 201.04) = 0.8678

Weight of Debt = 1 - 0.8678 = 0.1322

As per CAPM Model,

Cost of Equity = Rf + Beta(Rm - Rf)

Cost of Equity = 0.06 + 1.20(0.12 - 0.06)

Cost of Equity = 13.20%

After-tax cost of debt = (1 - t)Cd

After-tax cost of debt = (1 - 0.36)(0.108)

After-tax cost of debt = 6.91%

WACC = (0.1322)(0.0691) + (0.8678)(0.1320)

WACC = 12.37%